# Offline-to-Online Multi-Agent Reinforcement Learning with Offline Value Function Memory and Sequential Exploration

### v1

@@ 5.3. Ablation Study
In this section, we conduct additional experiments to demonstrate the effectiveness of both OVM and Sequential Exploration. In the previous section, we established the superiority of the OVMSE approach. Through ablation studies, we aim to answer two key questions:
@@
We perform these ablation studies on the 2s3z medium task, 6h_vs_8z medium task, and medium replay tasks. As shown in Figure 4(a), (b), and (c), we set the mixing ratio to 0.0 for all tasks except for the 6h_vs_8z medium task, where a mixing ratio of 0.1 is applied. Additionally, we conduct a specific ablation study on the 2s3z medium task to examine the effect of reusing offline data during the online phase. As illustrated in Figure 4, we report the mean test returns as a function of environment steps, as this metric more effectively captures performance drops during the initial stages of online learning.
OVM is effective. Our results demonstrate that OVM plays a crucial role in O2O MARL. In the 2s3z medium task, using SE alone without OVM results in a more significant performance drop during the initial stage of the online phase. Additionally, using OVM alone results in a much smaller performance drop compared to the action proposal approach Hu et al. (2024a). In the 6h_vs_8z medium replay task, OVMSE significantly outperforms SE alone, due to the preservation of offline knowledge by OVM. Similarly, in the 6h_vs_8z medium task, OVMSE outperforms SE, with SE requiring 2 million steps to catch up.
@@
OVMSE has minimal reliance on offline data. As shown in Figure 4(d), MACal-QL experiences a much larger performance drop when the mixing ratio is set to $0.0$ compared to $0.5$. This occurs because reusing offline data during online training helps reduce distribution shift. However, over-reliance on offline data can lead to slow online fine-tuning, especially when the offline data is of poor quality. In contrast, OVMSE with a mixing ratio of $0.0$ achieves fast and stable online fine-tuning during the online phase, demonstrating that OVMSE is not sensitive to data distribution change and well preserves the learned Q-values.
@@ 5.4. Discussions
In this section, we discuss the role of offline pre-trained knowledge in O2O MARL. Offline knowledge is beneficial in several ways. First, it helps guide online exploration much more efficiently than training from scratch. It also provides a strong initial policy. Furthermore, the offline pre-trained knowledge can facilitate faster online learning and lead to better performance, as clearly demonstrated in the 6h_vs_8z medium replay task shown in Figure 4 (b), where OVMSE outperforms SE by a large margin, thanks to its offline knowledge.
However, we also observe that offline knowledge can hinder online learning in some cases. This issue likely arises because current offline training algorithms are not designed for the online phase. A good offline policy requires some degree of over-fitting to the offline datasets (e.g., achieved using pessimism or behavior cloning). However, for O2O MARL, this over-fitting can cause online learning to require many samples to correct or even result in the agent sticking to the offline policy, thereby preventing improvement. This observation motivated the introduction of the memory coefficient $\lambda_{\text{memory}}$ and its corresponding annealing schedule, which allow OVMSE to gracefully adjust offline knowledge to the level of online returns.
@@ 6. CONCLUSIONS
In this work, we investigate the O2O MARL. We identify two major challenges for O2O MARL: (i) During the initial stage of online learning, distributional shift can degrade the offline pre-trained Q-values, leading to unlearning of previously acquired knowledge; (ii) Inefficient exploration in the exponentially large joint state-action space can slow down online fine-tuning. To address these challenges, we introduce Offline Value Function Memory (OVM) to preserve offline knowledge and a decentralized sequential exploration strategy for efficient online exploration. Our approach, OVMSE, enables stable and fast online fine-tuning and leads to improved fine-tuned results. Experiments across various tasks in SMAC demonstrate the superior performance of OVMSE.

### v2

@@ 5.3. Ablation Study
In this section, we conduct additional experiments to demonstrate the effectiveness of both OVM and Sequential Exploration. Through ablation studies, we aim to answer three key questions:
@@
  3. (iii)
    Does OVMSE rely on reusing offline data during online phase?
We perform these ablation studies on the 2s3z medium task, 6h_vs_8z medium and medium replay tasks. As shown in Figure 4(a), (b), and (c), we set the mixing ratio to 0.0 for all tasks except for the 6h_vs_8z medium task, where a mixing ratio of 0.1 is applied. Additionally, we conduct a specific ablation study on the 2s3z medium task to examine the effect of reusing offline data during the online phase. As illustrated in Figure 4, we report the mean test returns as a function of environment steps, as this metric more effectively captures performance drops during the initial stages of online learning.
OVM is effective. Our results demonstrate that OVM plays a crucial role in O2O MARL. In the 2s3z medium task, using SE alone without OVM results in a more significant performance drop during the initial stage of the online phase. Additionally, using OVM alone results in a much smaller performance drop compared to the action proposal approach ImitationBoostrap. In the 6h_vs_8z medium replay task, OVMSE significantly outperforms SE alone, due to the preservation of offline knowledge by OVM. Similarly, in the 6h_vs_8z medium task, OVMSE outperforms SE, with SE requiring 2 million steps to catch up.
@@
OVMSE has minimal reliance on offline data during online training. We conduct an ablation study on 2s3z medium to examine the effect of reusing offline data during online training for OVMSE. We report the test returns at 200k environment steps (instead of the final fine-tuned test returns, as different mixing ratio values all achieve an optimal final test return). As shown in Table 2, OVMSE with a mixing ratio of 0.0 and 0.1 achieves fast and stable online fine-tuning during the online phase, demonstrating that OVMSE is not sensitive to data distribution shift and effectively preserves offline learned Q-values. This highlights a clear advantage of OVMSE, as over-reliance on offline data can lead to slow online fine-tuning, especially when the offline data is of poor quality.
@@ 5.4. Discussions
In this section, we discuss the role of offline pre-trained knowledge in O2O MARL. Offline knowledge is beneficial in several ways. First, it helps guide online exploration much more efficiently than training from scratch. It also provides a strong initial policy. Furthermore, the offline pre-trained knowledge can facilitate faster online learning and lead to better performance, as clearly demonstrated in the 6h_vs_8z medium replay task shown in Figure 4 (b), where OVMSE outperforms SE by a large margin, thanks to its offline knowledge. However, we also observe that offline knowledge can hinder online learning in some cases. This issue likely arises because current offline training algorithms are not designed for the online phase. A good offline policy requires some degree of over-fitting to the offline datasets. However, for O2O MARL, this over-fitting can cause online learning to require many samples to correct or even result in the agent sticking to the offline policy, thereby preventing improvement. This observation motivated the introduction of the memory coefficient $\lambda_{\text{memory}}$ and its corresponding annealing schedule, which allow OVMSE to gracefully adjust offline knowledge to the level of online returns.
@@ 6. CONCLUSIONS
In this work, we investigate the O2O MARL. We identify two major challenges for O2O MARL: (i) During the initial stage of online learning, distributional shift can degrade the offline pre-trained Q-values, leading to unlearning of previously acquired knowledge; (ii) Inefficient exploration in the exponentially large joint state-action space can slow down online fine-tuning. To address these challenges, we introduce Offline Value Function Memory (OVM) to preserve offline knowledge and decentralized sequential exploration strategy for efficient online exploration. Our approach, OVMSE, enables stable and fast online fine-tuning and leads to improved fine-tuned results. Experiments across various tasks in SMAC demonstrate the superior performance of OVMSE.
{acks}
This work is supported by the National Natural Science Foundation of China Grants 52450016 and 52494974.
Supplementary Materials
Appendix A Detailed Hyperparameters
In this section, we list in detail the hyperparameters used for OVMSE and all baseline algorithms.
A.1. Shared Hyperparameters
For all methods, we aim to keep the majority of hyperparameters consistent with the PYMARL2 implementation hu2023rethinkingtheimplementation, as these hyperparameters have been carefully fine-tuned. Table 3 outlines the differences in hyperparameters compared to PYMARL2, along with the coefficients related to offline training. These hyperparameters are shared across all algorithms, with the exception of QMIX training from scratch. For QMIX, we retain all hyperparameters identical to those in PYMARL2, except for Batch_Size_Run and Target_Update_Interval, which we adjust to match the other baseline methods for a fair comparison.
Table 3. Shared hyperparameter for different tasks.
Hyperparameters Tasks
2s3z 3s5z 5m_vs_6m 6h_vs_8z
Batch_Size_Run 2 2 8 8
Target_Update_Interval 50 50 200 200
Online_Gradient_Norm_Clip 1 0.5 10 10
Offline_CQL_Weight 0.05 0.25 50 20
A.2. Hyperparameters for Online Fine-tuning
For online fine-tuning, we mainly adjusted the online CQL weight, the mixing ratio and the exploration coefficient.
Mixing Ratio. For all algorithms except QMIX, we perform a search on the mixing ratio from the set $\{0.0,0.1,0.3,0.5\}$ and choose the best mixing ratio for each algorithm. The results are shown in Table 4.
Exploration Coefficient. For all algorithms except QMIX, we use $\epsilon_{\text{start}}=1$ and $\epsilon_{\text{end}}=0.05$ in all O2O tasks. For QMIX, the hyperparameters for $\epsilon$-greedy exploration are the same for all tasks as in PYMARL2 hu2023rethinkingtheimplementation.
Online CQL Weight. For MACQL and MACal-QL, we set the online CQL weight equal to the offline CQL weight for all tasks. OVMSE uses the same online CQL weight for 2s3z and the same online CQL weight as Switch CQL for the remaining tasks. The detailed settings are shown in Table 5.
Table 4. The mixing ratio of different algorithms for all datasets.
Datasets OVMSE Switch CQL MACal-QL MACQL
2s3z-m 0.0 N/A 0.5 0.3
2s3z-mr 0.0 N/A 0.3 0.3
3s5z-m 0.1 0.5 0.5 0.3
3s5z-mr 0.1 0.5 0.3 0.3
5m_vs_6m-m 0.1 0.1 0.5 0.5
5m_vs_6m-mr 0.0 0.3 0.5 0.5
6h_vs_8z-m 0.1 0.1 0.5 0.5
6h_vs_8z-mr 0.0 0.1 0.5 0.5
-m means medium, -mr means medium-replay.
Table 5. The online CQL weight of different algorithms for all datasets.
Datasets OVMSE Switch CQL MACal-QL MACQL
2s3z-m 0.05 N/A 0.05 0.05
2s3z-mr 0.05 N/A 0.05 0.05
3s5z-m 0.0 0.0 0.25 0.25
3s5z-mr 0.0 0.0 0.25 0.25
5m_vs_6m-m 0.0 0.0 50 50
5m_vs_6m-mr 0.0 0.0 50 50
6h_vs_8z-m 0.0 0.0 20 20
6h_vs_8z-mr 0.0 0.0 20 20
-m means medium, -mr means medium-replay.
A.3. Specific Hyperparameters for OVMSE
Featuring Offline Value Memory, the memory coefficient plays a crucial role in the performance of OVMSE. Table 6 provides the detailed settings of $\lambda_{\text{memory}}$ and its annealing schedule across different datasets. Additionally, Table 7 lists the exploration hyperparameters used by OVMSE with Sequential Exploration (SE).
Table 6. The memory coefficient annealing schedule for all datasets.
Datasets $\lambda_{\text{memory\_start}}$ $\lambda_{\text{memory\_end}}$ Annealing Duration $T$
2s3z-m 1.0 1.0 N/A
2s3z-mr 1.0 1.0 N/A
3s5z-m 1.0 1.0 N/A
3s5z-mr 1.0 1.0 N/A
5m_vs_6m-m 1.0 0.1 200k
5m_vs_6m-mr 1.0 0.05 200k
6h_vs_8z-m 1.0 0.05 100k
6h_vs_8z-mr 1.0 0.05 100k
-m means medium, -mr means medium-replay.
Table 7. Exploration hyperparameters for OVMSE with Sequential Exploration (SE).
Hyperparameters Value
Exploration Strategy Sequential Exploration (SE)
$\epsilon_{\text{start}}$ 1.0
$\epsilon_{\text{end}}$ 0.05
$\epsilon$ Annealing Duration 100k timesteps
A.4. Offline CQL Algorithm Choice
The CQL-based offline MARL algorithm we choose for 2s3z and 3s5z is CQL, while we use CFCQL CFCQL for 5m_vs_6m and 6h_vs_8z due to its better offline performance. All algorithms use the same CQL loss for the same tasks.
Appendix B Details on Data Collection
Figure 5. The distribution of trajectories with different rewards in the 3s5z medium and 3s5z medium-replay datasets. The horizontal axis represents the return of each trajectory, and the vertical axis represents the percentage of trajectories corresponding to that return.
Since an offline dataset for the SMAC grandmasterstarcraft2 3s5z task is currently unavailable, we created our own medium and medium-replay datasets. First, we trained a multi-agent policy from scratch to a medium level by running QMIX, implemented by PYMARL2, for $100,000$ environment interaction steps. For the medium dataset, we executed the trained medium-level policy in the environment and collected $5,000$ episodes. For the medium-replay dataset, we saved the replay buffer from the QMIX agent at the final training step (i.e., the $100,000$th step). Details of the datasets are provided in Table 8, and the distribution of the datasets are shown in Figure 5.
Table 8. Primary information about our own dataset.
Datasets Size Min Return Average Return Max Return
3s5z medium 5,000 Episodes 2.57 11.15 20.03
3s5z medium-replay 1,908 Episodes 2.29 7.05 20.16